Mr Smith has 15 3/4 yards of plastic strips for making bracelets. she needs 1/3 yard for each bracelet. how many whole bracelets can she make?
Answer by rfer(16322)

(63/4)/(1/3)=
63/4*3/1=189/4
189/4=47 plus 1/4 of one